# Connecting GRBs from Binary Neutron Star Mergers to Nuclear Properties of Neutron Stars

### v1

@@ Abstract
The fate of the binary neutron star (NS) merger remnants hinges sensitively upon the NS equation of state and the threshold mass, $M_{\rm ls}$, that separates a long-lived from a short-lived NS remnant. The nature of the electromagnetic counterparts is also influenced by the remnant type, particularly in determining whether a gamma-ray burst from a compact binary merger (cbGRB) is of short or long duration. We propose a novel approach to probe the threshold mass by linking it to the estimated observed ratio of long to short cbGRBs. We find that current observations broadly favour a relatively high value for this transition, $M_{\rm ls}\simeq 1.3M_{\rm TOV}$, for which $M_{\rm TOV}\lesssim 2.6\,M_{\odot}$, consistent with numerical simulations, as also shown here. Our results disfavour nuclear physics scenarios that would lead to catastrophic pressure loss at a few times nuclear density and temperatures of tens of MeV, leading to a rapid gravitational collapse of binaries with total mass $M\lesssim 1.3M_{\rm TOV}$. Future individual gravitational wave events with on-axis cbGRBs can further bound $M_{\rm ls}$.
@@
A recent analysis by Rastinejad et al. [26] indicates that lbGRBs are accompanied by bright red kilonovae, consistent with an origin in massive BH disks, while sbGRBs are linked to luminous blue kilonovae. Gottlieb et al. [25] showed that the brightness and color of kilonovae associated with sbGRBs cannot be accounted for by BH-powered jets alone. They concluded that long-lived remnant NSs are likely the central engines of sbGRBs, whereas low BH disks rather power fainter lbGRBs. In fact, they pointed out that GW170817 was likely an lbGRB rather than an sbGRB.
The framework of Gottlieb et al. [25], which connects diverse GRB classes from BNS mergers with kilonova properties, remnant types, and post-merger accretion disk masses, also enables the extraction of essential information about the NS EoS. The observed relative rates of various GRB types can constrain the parameter space of viable EoS models [27]. This approach has been previously utilized by Piro et al. [28], Sarin et al. [29], Guglielmi et al. [30], who constrained the EoS based on the fraction of short GRBs with plateaus, under the assumption that these plateaus indicate a remnant NS passing through a magnetar phase either as a stable NS, or a NS remnant collapsing to a BH after losing its centrifugal support.
Building on the recent unification model by Gottlieb et al. [24, 25] and current observational constraints on the relative rates of lbGRBs to sbGRBs, we provide a novel approach to constraining nuclear properties of NS matter, with particular sensitivity to the threshold NS mass setting the transition between a long- and short-lived NS remnant. The paper is organized as follows. In Section II, we describe the key model ingredients of our analysis: the intrinsic distribution of NS masses, the binary NS merger remnant types, and the modeling of the disk mass remnant as a function of the total BNS mass and their mass ratio. In Section IIIA, we present the results of our population study, focusing on relative fractions of the merger remnants and their surrounding disks. These are transformed in Section IIIB into theoretical predictions on the relative rates of lbGRBs to sbGRBs according to the model of Gottlieb et al. [24, 25]. We compare these predictions with observations, to constrain the threshold mass between a long-lived and a short-lived NS remnant. In Section IIIC, we compare our results to numerical simulations and find a remarkable consistency. We conclude and highlight the implications of our results for the nuclear properties of NS matter in Section IV.
@@ II.1 Intrinsic distributions of NS masses
The mass distribution of the BNS merger remnants closely hinges upon the mass distribution of the NS binary components. Galactic NS star systems are known to have a distribution tightly peaked around $1.35M_{\odot}$, but the mass distribution of all known pulsars is now known to be broader [31], with the current range spanning from $1.17M_{\odot}$ to $2.35\pm 0.17M_{\odot}$ [32]. As the wealth of observations has been growing over the years combining measurements from different types of systems, it has become apparent that the NS mass distribution is bimodal [33, 34, 35, 36, 37], consistent with what was found in simulations of the explosions of massive stars [38].
@@
Lacking detailed predictions for the relative mass distributions of the two NS members of the binary (i.e. their mass ratio), and to avoid making assumptions that then influence the results, we choose to make no prior assumptions on the mass ratio: both NS masses are randomly generated from the same binary distribution, with the primary being the larger of any two randomly drawn values. The resulting distributions of the primary ($M_{1}$) and secondary ($M_{2}$) masses are displayed in the top panel of Fig. 1, while their mass ratio $q\equiv M_{2}/M_{1}$ is correspondingly shown in the bottom panel of the same figure. It can be seen that, with the no prior assumption made on the $q$ distribution, the natural probability favors comparable masses.
Figure 1: Top: Probability distributions of primary ($M_{1}$) and secondary ($M_{2}$) NS mass. Bottom: Corresponding probability distribution of the mass ratio $q$ under a ’no-prior assumption’ on their mass ratio.
II.2 Remnant type as a function of the EOS
Given two merging NSs of masses $M_{1}$ and $M_{2}$ (randomly drawn from the distributions above), we calculate the total remnant mass as $M_{\rm tot}=M_{1}+M_{2}$.
There are five possible types of remnants from a BNS merger [39]:
@@ II.2 Remnant type as a function of the EOS
    Very Long-lived Remnant Neutron Star (VLRNS): If $M_{\rm TOV}<M_{\rm tot}\leq M_{\rm L,sp}$, where $M_{\rm L,sp}$ is the maximum mass of a uniformly rotating NS secured by centrifugal and thermal effects that possibly collapses to a BH on a dynamical or secular timescale. We use the conservative lower-limit $M_{\rm L,sp}=1.15M_{\rm TOV}$.
@@
    Long-Lived Remnant Neutron Star (LRNS): If $M_{\rm L,sp}\lesssim M_{\rm tot}<aM_{\rm TOV}$. These are long-lived remnants that collapse on the viscous timescale. The mass $M_{\rm ls}\equiv aM_{\rm TOV}$ represents the threshold separating long- and short-lived remnants. The analysis presented in this work shows the potential of constraining it phenomenologically.
@@
    Short-lived Remnant Neutron Star (SRNS): If $M_{\rm ls}\leq M_{\rm tot}\leq M_{\rm thr}$, where $M_{\rm thr}$ is the threshold mass for prompt collapse to BH. These are short-lived remnants that collapse on the GW timescale.
@@ II.3 Disk Mass Remnant
The various best-fit parameters have the following numerical values: $a_{0}=-1.581$, $\delta a=-2.439$, $b_{0}=-0.538$, $\delta_{b}=-0.406$, $\beta=3.910$, $q_{\rm trans}=0.900$.
@@
III.1 The post-merger remnant population
@@
To connect the various elements laid out in the previous section, we perform Monte Carlo simulations to determine the properties of the remnant population from BNS mergers.
@@
Given the initial NS mass distribution, we randomly draw the binary NS masses from it. Within our simple, no-prior assumption model, which assumes that the masses of the two NSs in the binary are uncorrelated with each other, both the primary mass $M_{1}$ and the secondary mass $M_{2}$ are drawn from the same distribution. Given two random drawings, the larger of the two masses is assigned to be the primary and the other is the secondary.
@@
For any given merger remnant of mass $M_{\rm tot}$, we evaluate which of the five conditions listed in Sec. IIB is satisfied, and assign the remnant to the corresponding group. We perform this computation for a wide range of 52 EoSs taken from Table 2 of Ref. [41], spanning the current range of nuclear physics uncertainties, for which we use $M_{\rm TOV}$ as a proxy to classification.
@@
Figure 2: Remnant fractions as a function of the total mass of the binary merger remnant, where both $M_{1}$ and $M_{2}$ are independently drawn from the bimodal NS mass distribution. Here we adopted an intermediate, representative value $M_{\rm ls}=1.3M_{\rm TOV}$. A larger (smaller) value would increase (decrease) the relative abundance of LRNSs with respect to SRNSs.
@@
Figure 3: Post-BNS merger disk mass as a function of the NS EoS (parametrized via $M_{\rm TOV}$) and the mass ratio $q$ of the merging NSs. Here we used a representative value $M_{\rm ls}=1.3M_{\odot}$ for the LRNS to SRNS transition. A larger value would shift some remnants from the SRNS+pcBH sample to that of LRNS in the bottom plots, and vice versa for a smaller value.
@@
Fig. 2 shows the contribution of each remnant group, as a function of $M_{\rm TOV}$. The overall trends are as expected: the larger $M_{\rm TOV}$, the higher the relative fraction of stable NS remnants, and the lower the fraction of prompt-collapse BHs. While the fractions of SNSs, VLRNSs, and LRNSs follow a relatively smooth trend (other than for statistical fluctuations), the fractions of SRNSs and pcBHs display relatively large fluctuations deviating from the overall trend. This stems from the fact that, while we have ordered the EoS by their $M_{\rm TOV}$, the corresponding values of pcBH do not strictly follow the same trends, i.e. some EoS may have a larger $M_{\rm TOV}$ but a smaller pcBH, hence resulting in non-strictly monotonic trends for the corresponding SRNS and pcBH remnant fractions when plotted versus $M_{\rm TOV}$.
@@
For each BNS merger, we estimate the disk mass using the fit in Eq. 3. We separate the mass distributions for each remnant type but combine the contributions of the SRNSs and the pcBHs since they yield similar outcomes in our analysis. The disk masses are displayed as a function of both $M_{\rm TOV}$ and $q$ in Fig. 3. The dependence of the disk mass on the mass ratio is evident for all the remnant types, but especially pronounced for the SRNS+pcBH population, where the disk mass can vary from several tens of solar masses for high unequal mass ratios to negligible disk mass in the limit of equal mass ratios. Stable NS remnants have a generally large disk mass, from about $\sim 0.5M_{\odot}$ for a small mass ratio down to $\sim 0.15-0.2M_{\odot}$ (with the precise value depending on the EoS) for equal mass ratios. The disk mass for the VLRNS population bears similarities to the stable NS while the LRNS case is more similar to the SRNS+pcBH population, albeit with a gentler dependence on the mass ratio. Note that the structure in the disk mass, especially pronounced for the populations of LRNS and SRNS+pcBH, comes from the fact that $M_{\rm thr}$ is not directly proportional to $M_{\rm TOV}$.
@@ III.2 Long and Short GRBs from BNS mergers
We connect the properties of the compact remnant types and disk-mass distributions with the findings of Gottlieb et al. [24] and the rate information to date of the lbGRB/sbGRB fraction. From the compilation of Fong et al. [47], this fraction in the $z<0.3$ sample can be estimated at about 40-50%, using GRB 060614 [48], GRB 211211A [17] and GRB230307A [49] to represent the lbGRB share. We note that this is a rough estimate and possibly subject to change with more future data. A more comprehensive analysis extending the sample to $z\lesssim 0.5$ suggests that the long-to-short ratio might be even as high as $\sim 1$ (A. Levan, private communication). However, despite current uncertainties in the data, our goal here is to show the constraining power for the nuclear properties of NSs of connecting theoretical ideas for the compact binary GRB phenomenology with the observed population. While in this work we derive constraints based on the current data, the analysis can be refined with a larger wealth of observations in the future.
@@
    Millisecond SNSs and VLRNSs are disfavoured as GRB progenitors since their immense rotational energy is notably absent in GRB radio afterglows [51, 52, 53, 54] and optical kilonova signals. Numerical simulations of BNS mergers also generally suggest that a hypermassive NS phase is likely necessary to launch a successful GRB jet (see e.g. [55, 56, 57], but also note [58] for a successful outcome). While we adopt this as our fiducial model, however we also explore how our results vary if this assumption is released and magnetars associated with VLRNSs make up a sizeable fraction of cbGRBs.
@@
    The outcome of mergers resulting in SRNSs or prompt BH formation depends on the disk mass. When the disk mass is high, $M_{d}\gtrsim 0.1\,M_{\odot}$, the BH will inevitably produce a long-lived jet, leading to a detectable lbGRB. However, if the disk mass is $M_{d}\lesssim 10^{-2}\,M_{\odot}$, then given the magnetic flux on the BH, proportional to $B^{2}\sim M_{d}$, is reduced, the jet would be too faint to detect.
@@
    LRNSs power the standard sbGRBs, where the NS remnant lifetime roughly determines the GRB duration. We note that LRNSs are predicted to form massive accretion disks [e.g., 7]. This implies that in some cases, the collapse of the LRNSs into a BH could extend the emission, potentially producing an lbGRB. However, the classification of the GRB would depend on the relative power of the NS-driven and BH-driven jets [see discussion in 25]. While some LRNSs may contribute to the lbGRB population, for the following analysis we assume that all LRNSs contribute exclusively to the sbGRB population.
@@
To match the observed GRB populations, the rate of LRNSs must be comparable to that of SRNSs and pcBHs with massive disks, which produce detectable lbGRBs.
@@
Fig. 4 depicts the ratio of LRNSs to (SRNSs + pcBHs) with disk mass $\geq 0.1M_{\odot}$ as a function of $M_{\rm TOV}$, and for five different choices of the transition mass $M_{\rm ls}$ in a broad range of values between 1.2-1.4 $M_{\rm TOV}$ The grey band guides the eye to the current estimated rates, for which we have adopted a broad conservative range of $50\%-100\%$. Taking these rates at face value, $M_{\rm ls}=1.2M_{\rm TOV}$ is largely inconsistent with the data, for any choice of the NS EoS. A transition at $M_{\rm ls}=1.25M_{\rm TOV}$ would be compatible with the data but only for extremely stiff EoS, with $M_{\rm TOV}\gtrsim 2.7M_{\odot}$, which are disfavoured by GW170817 observations [4, 9, 10]. On the other hand, a transition around the higher value $M_{\rm ls}=1.3M_{\rm TOV}$ is consistent with the data for a range of acceptable EoS with $M_{\rm TOV}\lesssim 2.6M_{\odot}$. Similarly, a transition at $M_{\rm ls}=1.35M_{\rm TOV}$ is acceptable within a broad range of viable EoSs with $2.1M_{\odot}\lesssim M_{\rm TOV}\lesssim 2.4M_{\odot}$, while the very high value $M_{\rm ls}=1.4M_{\rm TOV}$ can be accommodated, but only within the softer EoS range with $M_{\rm TOV}\lesssim 2.2M_{\odot}$. Therefore, we conclude that the transition between the long- and short-lived NS remnants likely occurs in the range $\sim[1.3-1.4]M_{\rm TOV}$.
@@
Figure 4: The ratio between sbGRBs and lbGRBs, where the former are identified with the LRNSs and the latter with the combined populations of SRNSs and pcBH with a disk mass larger than a minimum value. Five different limits are considered for the transition mass $M_{\rm ls}$ between a long-lived and a short-lived NS remnant. The grey band denotes the current observational constraints.
@@
We note that a premise of our analysis is that the contribution to lbGRBs from NS–BH mergers is negligible compared to NS–NS mergers. This is motivated by the fact that numerical relativity simulations of NS–BH mergers [59] show, for a variety of mass ratios and BH spin, that the disk mass is generally small. Massive disks can only be obtained for less massive BHs or very rapidly spinning BHs [^1] Both NS–BH events detected to date are in the regime of high-mass ratio and low spin [60]. Thus, if representative of the population, detectable EM counterparts are likely to be uncommon.. Hence, even if NS–BH mergers were to contribute to lbGRBs, they would likely be too dim to detect.

### v2

@@ Abstract
The fate of the binary neutron star (NS) merger remnants hinges sensitively upon the NS equation of state and the threshold mass, $M_{\rm ls}$, that separates a long-lived from a short-lived NS remnant. The nature of the electromagnetic counterparts is also influenced by the remnant type, particularly in determining whether a gamma-ray burst from a compact binary merger (cbGRB) is of short or long duration. We propose a novel approach to probe $M_{\rm ls}$ by linking it to the estimated observed ratio of long to short cbGRBs. We find that current observations broadly favour a relatively high value for this transition, $M_{\rm ls}\simeq 1.3M_{\rm TOV}$, for which $M_{\rm TOV}\lesssim 2.6\,M_{\odot}$, consistent with numerical simulations, as also shown here. Our results disfavour nuclear physics scenarios that would lead to catastrophic pressure loss at a few times nuclear density and temperatures of tens of MeV, leading to a rapid gravitational collapse of binaries with total mass $M\lesssim 1.3M_{\rm TOV}$. Future individual gravitational wave events with on-axis cbGRBs can further bound $M_{\rm ls}$.
@@
A recent analysis by Rastinejad et al. [26] indicates that lbGRBs are accompanied by bright red kilonovae, consistent with an origin in massive BH disks, while sbGRBs are linked to luminous blue kilonovae. Gottlieb et al. [25] showed that the brightness and color of kilonovae associated with sbGRBs cannot be accounted for by BH-powered jets alone. They concluded that long-lived remnant NSs are likely the central engines of sbGRBs, whereas low-mass BH disks rather power fainter lbGRBs. In fact, they pointed out that GW170817 was likely an lbGRB rather than an sbGRB.
The framework of Gottlieb et al. [25], which connects diverse GRB classes from BNS mergers with kilonova properties, remnant types, and post-merger accretion disk masses, also enables the extraction of essential information about the NS EoS. The connection between remnant types from NS-NS mergers, short GRBs, and the NS EoS has long been noted in the literature. Belczynski et al. [27] and Fryer et al. [28] pointed out that $M_{\rm TOV}$ can be constrained under the assumptions that short GRBs from NS-NS mergers require a BH remnant. Salafia et al. [29] discussed a method to constrain the NS EoS and mass distribution using the fraction of GW NS-NS mergers with an associated jet, with the latter assumed to require a BH remnant and a minimum disk mass of $\sim 10^{-3}M_{\odot}$. The statistical approach in terms of incidence of remnant types has also been previously utilized by Piro et al. [30], Sarin et al. [31], Guglielmi et al. [32], who constrained the EoS based on the fraction of short GRBs with plateaus, under the assumption that these plateaus indicate a remnant NS passing through a magnetar phase either as a stable NS, or a NS remnant collapsing to a BH after losing its centrifugal support.
Here, building on the recent unification model by Gottlieb et al. [24, 25] and current observational constraints on the relative rates of lbGRBs to sbGRBs, we provide a novel approach to constraining nuclear properties of NS matter which has particular sensitivity to the threshold NS mass setting the transition between a long- and short-lived NS remnant. The paper is organized as follows. In Section II, we describe the key model ingredients of our analysis: the intrinsic distribution of NS masses, the binary NS merger remnant types, and the modeling of the disk mass remnant as a function of the total BNS mass and their mass ratio. In Section IIIA, we present the results of our population study, focusing on relative fractions of the merger remnants and their surrounding disks. These are transformed in Section IIIB into theoretical predictions on the relative rates of lbGRBs to sbGRBs according to the model of Gottlieb et al. [24, 25]. We compare these predictions with observations, to constrain the threshold mass between a long-lived and a short-lived NS remnant. In Section IIIC, we compare our results to numerical simulations and find a remarkable consistency. We conclude and highlight the implications of our results for the nuclear properties of NS matter in Section IV.
@@ II.1 Intrinsic distributions of NS masses
The mass distribution of the BNS merger remnants closely hinges upon the mass distribution of the NS binary components. Galactic NS star systems are known to have a distribution tightly peaked around $1.35M_{\odot}$, but the mass distribution of all known pulsars is now known to be broader [33], with the current range spanning from $1.17M_{\odot}$ to $2.35\pm 0.17M_{\odot}$ [34]. As the wealth of observations has been growing over the years combining measurements from different types of systems, it has become apparent that the NS mass distribution is bimodal [35, 36, 37, 38, 39, 40], consistent with what was found in simulations of the explosions of massive stars [41].
Figure 1: Top: Probability distributions of primary ($M_{1}$) and secondary ($M_{2}$) NS mass. Bottom: Corresponding probability distribution of the mass ratio $q$ under a random pairing assumption on their mass ratio.
@@
Lacking detailed predictions for the relative mass distributions of the two NS members of the binary (i.e. their mass ratio), and to avoid making assumptions that then influence the results, we choose to make no prior assumptions on the mass ratio: both NS masses are randomly generated from the same binary distribution  [^1] A note of caution is that, for main sequence stars in binaries, observations show that random pairing is not supported [42]. For NS in binaries, more data will be needed to draw definite conclusions., with the primary being the larger of any two randomly drawn values. The resulting distributions of the primary ($M_{1}$) and secondary ($M_{2}$) masses are displayed in the top panel of Fig. 1, while their mass ratio $q\equiv M_{2}/M_{1}$ is correspondingly shown in the bottom panel of the same figure. It can be seen that, with the random pairing assumption made on the $q$ distribution, the natural probability favors comparable masses.
This is the NS mass model we designate as our fiducial choice. In the appendix, we will explore how varying this model in different ways affects our results.
II.2 Merger outcome as a function of the EOS
Given two merging NSs of masses $M_{1}$ and $M_{2}$ (randomly drawn from the distributions above), we calculate the total mass of the binary when the stars are at infinite separation as $M_{\rm tot}=M_{1}+M_{2}$. We note that this should not be intended as the baryonic or gravitational merger remnant often used in the literature (e.g. [30, 10, 29]).
There are five possible outcomes from a BNS merger [43]:
@@ II.2 Merger outcome as a function of the EOS
    Very Long-lived Neutron Star (VLNS): If $M_{\rm TOV}<M_{\rm tot}\leq M_{\rm L,sp}$, where $M_{\rm L,sp}$ is the maximum mass of a uniformly rotating NS secured by centrifugal and thermal effects that possibly collapses to a BH on a dynamical or secular timescale. While it has been shown that rigid rotation can support stars with mass up to 1.2 $M_{\rm TOV}$ [44], here we use the lower-limit $M_{\rm L,sp}=1.15M_{\rm TOV}$. This is a conservative limit for the results presented in Sec. III, as quantitatively shown in the appendix.
@@
    Long-Lived Neutron Star (LLNS): If $M_{\rm L,sp}\lesssim M_{\rm tot}<aM_{\rm TOV}$. These are long-lived objects that collapse on the viscous timescale. The mass $M_{\rm ls}\equiv aM_{\rm TOV}$ represents the threshold separating long- and short-lived remnants. The analysis presented in this work shows the potential of constraining it phenomenologically.
@@
    Short-lived Neutron Star (SLNS): If $M_{\rm ls}\leq M_{\rm tot}\leq M_{\rm thr}$, where $M_{\rm thr}$ is the threshold mass for prompt collapse to BH. These are short-lived objects that collapse on the GW timescale.
@@
While using Eq. 1 and 2 as part of our fiducial model, in the appendix we will quantitatively explore the effect of variations from this adopted fit.
@@ II.3 Disk Mass Remnant
The various best-fit parameters have the following numerical values: $a_{0}=-1.581$, $\delta a=-2.439$, $b_{0}=-0.538$, $\delta_{b}=-0.406$, $c=0.953$, $d=0.0417$, $\beta=3.910$, $q_{\rm trans}=0.900$.
This disk fit is a component of our fiducial model. However, in the appendix we will explore how varying it in different ways impacts our results.
@@
III.1 The post-merger object population
@@
To connect the various elements laid out in the previous section, we perform Monte Carlo simulations to determine the frequency of the various outcomes and the corresponding disk mass distributions from BNS mergers.
@@
Given the initial NS mass distribution, we randomly draw the binary NS masses from it. Within our simple, random pairing assumption model, which assumes that the masses of the two NSs in the binary are uncorrelated with each other, both the primary mass $M_{1}$ and the secondary mass $M_{2}$ are drawn from the same distribution. Given two random drawings, the larger of the two masses is assigned to be the primary and the other is the secondary.
@@
For any binary of total mass $M_{\rm tot}$, we evaluate which of the five conditions listed in Sec. IIB is satisfied, and assign the outcome to the corresponding group. We perform this computation for a wide range of 52 EoSs taken from Table 2 of Ref. [46], spanning the current range of nuclear physics uncertainties, for which we use $M_{\rm TOV}$ as a proxy to classification.
@@
Figure 2: Remnant fractions as a function of the total mass of the binary merger remnant, where both $M_{1}$ and $M_{2}$ are independently drawn from the bimodal NS mass distribution. Here we adopted an intermediate, representative value $M_{\rm ls}=1.3M_{\rm TOV}$. A larger (smaller) value would increase (decrease) the relative abundance of LLNSs with respect to SLNSs.
@@
Figure 3: Post-BNS merger disk mass as a function of the NS EoS (parametrized via $M_{\rm TOV}$), the mass ratio $q$ of the merging NSs, and the total merger mass $M_{\rm tot}$. Here we used a representative value $M_{\rm ls}=1.3M_{\odot}$ for the LLNS to SLNS transition. A larger value would shift some remnants from the SLNS+pcBH sample to that of LLNS in the bottom plots, and vice versa for a smaller value.
@@
Fig. 2 shows the contribution of each group, as a function of $M_{\rm TOV}$. The overall trends are as expected: the larger $M_{\rm TOV}$, the higher the relative fraction of stable NSs, and the lower the fraction of prompt-collapse BHs. While the fractions of SNSs, VLNSs, and LLNSs follow a relatively smooth trend (other than for statistical fluctuations), the fractions of SLNSs and pcBHs display relatively large fluctuations deviating from the overall trend. This stems from the fact that, while we have ordered the EoS by their $M_{\rm TOV}$, the corresponding values of pcBH do not strictly follow the same trends, i.e. some EoS may have a larger $M_{\rm TOV}$ but a smaller pcBH, hence resulting in non-strictly monotonic trends for the corresponding SLNS and pcBH remnant fractions when plotted versus $M_{\rm TOV}$.
@@
For each BNS merger in our MC simulations, we estimate the disk mass using the fit in Eq. 3. We separate the mass distributions for each outcome but combine the contributions of the SLNSs and the pcBHs since they yield similar outcomes in our analysis. The disk masses are displayed as a function of $M_{\rm TOV}$, $q$ and $M_{\rm tot}$ in Fig. 3. The dependence of the disk mass on the mass ratio is evident for all the cases, but especially pronounced for the SLNS+pcBH population, where the disk mass can vary from several tens of solar masses for high unequal mass ratios to negligible disk mass in the limit of equal mass ratios. Stable NSs have a generally large disk mass, from about $\sim 0.5M_{\odot}$ for a small mass ratio down to $\sim 0.15-0.2M_{\odot}$ (with the precise value depending on the EoS) for equal mass ratios. The disk mass for the VLNS population bears similarities to the stable NS while the LLNS case is more similar to the SLNS+pcBH population, albeit with a gentler dependence on the mass ratio. The scatter plots of Fig. 3 also illustrate the range of $M_{\rm tot}$, and the relative frequency, contributing to each type of merger outcome.
@@ III.2 Long and Short GRBs from BNS mergers
We connect the properties of the BNS merger outcomes with the findings of Gottlieb et al. [24] and the rate information to date of the lbGRB/sbGRB fraction. From the compilation of Fong et al. [52], this fraction in the $z<0.3$ sample can be estimated at about 40-50%, using GRB 060614 [53], GRB 211211A [17] and GRB230307A [54] to represent the lbGRB share. We note that this is a rough estimate and possibly subject to change with more future data. A more comprehensive analysis extending the sample to $z\lesssim 0.5$ suggests that the long-to-short ratio might be even as high as $\sim 1$ (A. Levan, private communication). However, despite current uncertainties in the data, our goal here is to show the constraining power for the nuclear properties of NSs of connecting theoretical ideas for the compact binary GRB phenomenology with the observed population. While in this work we derive constraints based on the current data, the analysis can be refined with a larger wealth of observations in the future.
@@
    Millisecond SNSs and VLNSs are disfavoured as GRB progenitors since their immense rotational energy is notably absent in GRB radio afterglows [56, 57, 58, 59] and optical kilonova signals. Numerical simulations of BNS mergers also generally suggest that a hypermassive NS phase is likely necessary to launch a successful GRB jet (see e.g. [60, 61, 62], but also note [63] for a successful outcome). While we adopt this as our fiducial model, however we also explore how our results vary if this assumption is released and magnetars associated with VLNSs make up a sizeable fraction of cbGRBs.
@@
    The outcome of mergers resulting in SLNSs or prompt BH formation depends on the disk mass. When the disk mass is high, $M_{d}\gtrsim 0.1\,M_{\odot}$, the BH will inevitably produce a long-lived jet, leading to a detectable lbGRB. However, if the disk mass is $M_{d}\lesssim 10^{-2}\,M_{\odot}$, then given the magnetic flux on the BH, proportional to $B^{2}\sim M_{d}$, is reduced, the jet would be too faint to detect.
@@
    LLNSs power the standard sbGRBs, where the NS remnant lifetime roughly determines the GRB duration. We note that LLNSs are predicted to form massive accretion disks [e.g., 7]. This implies that in some cases, the collapse of the LLNSs into a BH could extend the emission, potentially producing an lbGRB. However, the classification of the GRB would depend on the relative power of the NS-driven and BH-driven jets [see discussion in 25]. While some LLNSs may contribute to the lbGRB population, for the following analysis we assume that all LLNSs contribute exclusively to the sbGRB population.
@@
To match the observed GRB populations, the rate of LLNSs must be comparable to that of SLNSs and pcBHs with massive disks, which produce detectable lbGRBs.
@@
Fig. 4 depicts the ratio of LLNSs to (SLNSs + pcBHs) with disk mass $\geq 0.1M_{\odot}$ as a function of $M_{\rm TOV}$, and for five different choices of the transition mass $M_{\rm ls}$ in a broad range of values between 1.2-1.4 $M_{\rm TOV}$. We should however note that for 8/52 of the EoS explored here [^2] These are: APR ($M_{\rm TOV=}2.188M_{\odot}$), GMSR(H1) ($M_{\rm TOV=}2.289M_{\odot}$), GMSR(H3) ($M_{\rm TOV=}2.302M_{\odot}$), GMSR(H4) ($M_{\rm TOV=}2.336M_{\odot}$), GMSR(H5) ($M_{\rm TOV=}2.381M_{\odot}$), GMSR(DHSL69) ($M_{\rm TOV=}2.411M_{\odot}$), GMSR(DHSL59) ($M_{\rm TOV=}2.427M_{\odot}$), GMSR(H7) ($M_{\rm TOV=}2.512M_{\odot}$). , the condition $1.4M_{\rm TOV}<M_{\rm th}$ is not satisfied. While this does not affect the computed fractions (in this case the SLNS+pcBH population simply reduces to the pure pcBH in the Monte Carlo), however we should note that for those EoS the LLNS-SLNS transition must occur for $a_{\rm crit}<1.4$ and therefore the corresponding points for $a_{\rm crit}=1.4$ should not be interpreted as a constraint on this parameter. In those specific cases, the lbGRBs/sbGRBs ratio sensitivity is rather shifted to $M_{\rm th}$. The grey band in the figure guides the eye to the current estimated rates, for which we have adopted a broad conservative range of $50\%-100\%$. Taking these rates at face value, $M_{\rm ls}=1.2M_{\rm TOV}$ is largely inconsistent with the data, for any choice of the NS EoS. A transition at $M_{\rm ls}=1.25M_{\rm TOV}$ would be compatible with the data but only for extremely stiff EoS, with $M_{\rm TOV}\gtrsim 2.7M_{\odot}$, which are disfavoured by GW170817 observations [4, 9, 10]. On the other hand, a transition around the higher value $M_{\rm ls}=1.3M_{\rm TOV}$ is consistent with the data for a range of acceptable EoS with $M_{\rm TOV}\lesssim 2.6M_{\odot}$. Similarly, a transition at $M_{\rm ls}=1.35M_{\rm TOV}$ is acceptable within a broad range of viable EoSs with $2.1M_{\odot}\lesssim M_{\rm TOV}\lesssim 2.4M_{\odot}$, while the very high value $M_{\rm ls}=1.4M_{\rm TOV}$ can be accommodated, but only within the softer EoS range with $M_{\rm TOV}\lesssim 2.2M_{\odot}$. Therefore, we conclude that the transition between the long- and short-lived NS remnants likely occurs in the range $\sim[1.3-1.4]M_{\rm TOV}$.
@@
Figure 4: The ratio between sbGRBs and lbGRBs, where the former are identified with the LLNSs and the latter with the combined populations of SLNSs and pcBH with a disk mass larger than a minimum value. Five different limits are considered for the transition mass $M_{\rm ls}$ between a long-lived and a short-lived NS remnant. The grey band denotes the current observational constraints.
@@
We note that a premise of our analysis is that the contribution to lbGRBs from NS–BH mergers is negligible compared to NS–NS mergers. This is motivated by the fact that numerical relativity simulations of NS–BH mergers [64] show, for a variety of mass ratios and BH spin, that the disk mass is generally small. Massive disks can only be obtained for less massive BHs or very rapidly spinning BHs [^3] Both NS–BH events detected to date are in the regime of high-mass ratio and low spin [65]. Thus, if representative of the population, detectable EM counterparts are likely to be uncommon.. Hence, even if NS–BH mergers were to contribute to lbGRBs, they would likely be too dim to detect. However, more generally, note that any contamination to the observed sample of long cbGRBs by NS-BH mergers would mean that the fraction of long cbGRBs from NS-NS is smaller, hence strengthening our results.